OpenStudy (anonymous):

A tank contains 1000L of brine with 15kg of dissolved salt. Pure water enters the tank at the rate of 10L/min. The solution is kept thoroughly mixed and drains from the tank at the same rate. How much salt is in the tank (a) after t minutes and (b) after 20 minutes?

OpenStudy (amistre64):

No matter how much water enters the tank; there remains the same amount of salt as at the start of it all. It just becomes more diffuse.

OpenStudy (anonymous):

How do I set up the equation the answer is suppose to be 15e^-1/100 kg

OpenStudy (amistre64):

If its asking how diffuse the salt is; then its just a matter of calculating the amount of water added to the tank in the given time limit. 10 each minute after t minutes = 10t your addinging 10t liters of water that has no salt in it (0) to 1000 that has a dilution of 15kg 10t(0) + 1000(15) = (10t+1000) (15000 ..something else) so far

OpenStudy (amistre64):

lets adjust that to 20 minutes and see what we get :) 20(10) = 200 200(0) + 1000(15) = 1200(k) 15000 150 25 ------ = k = --- = ---- 1200 12 2 the final briney equates 25/2 in this case; or about 12.5 parts per liter

OpenStudy (amistre64):

to find "k" do this: 1000(15) 10(10(15)) --------- = ---------- = 150(t+10)^-1 10t+1000 10 (t+10) which is the discrete measurement; I believe "e" comes into the picture on a continuous measurement